OpenStudy (anonymous):

Please help me!!! If 832J of energy is required to raise the temperature of a sample of aluminum from 20.0°C to 97.0°C, what mass is the sample of aluminum? (The specific heat of aluminum is 0.90 J/(g × °C).) 0.10 g 10.0 g 12.0 g 57.7 g

OpenStudy (chmvijay):

Q= M *S *delta T M=Q /S*Delta T find the M Q is in joules energy S is specific heat capacity and delta T is temperature differeence

OpenStudy (anonymous):

So the answer is..?

OpenStudy (chmvijay):

u tell me i will tell its right or nooo:) cant u solve the above equation

OpenStudy (anonymous):

..Cannnttt u?

OpenStudy (chmvijay):

i cannn but u have to learn right :)

OpenStudy (anonymous):

Bla.. nvm

OpenStudy (chmvijay):

M= 832 / 0.9*77 tell me the answer common

OpenStudy (anonymous):

c

OpenStudy (chmvijay):

yaaa 12 g:)

OpenStudy (anonymous):

Thank u!!